OpenStudy (anonymous):

A car moving at 20.0m/s slows down at -1.50 m/s2 to a velocity of 10.0m/s. How far did the car go during the slow down?

OpenStudy (anonymous):

v^2-u^2=2as

OpenStudy (anonymous):

v is final velocity... u is initial velocity ... a is acc. and s is dist.

OpenStudy (shamim):

Or we can use V^2=u^2- 2as

OpenStudy (shamim):

Did u get it

OpenStudy (anonymous):

v^2=2as + u^2 take acc to be negative

OpenStudy (shamim):

Ya

OpenStudy (anonymous):

\[v ^{2}=2as+u^{2}\]